# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v3

@@ Appendix D Verlinde operator and topological defects: $\mathbb{C}$-linear category and fusion category
where $M$ and $\overline{M}$ are label of the chiral or antichiral descendant fields, corresponding to the Verma module. The most fundamental point in this representation is, by definition, $P_{a}$ is commutative with the CFT Hamiltonian $H_{CFT}$, i.e. $H_{CFT},P_{a}]=0$. In other words, the above form provides representations of conserved charges. Hence, the operators form a ring over $\mathbb{C}$. Moreover, by changing the basis by a linear transformation, one can introduce the following topological symmetry operators,
@@
by identifying $\times$ as the multiplication of the linear operators. By applying the modular $S$ transformation, the operator produces a defect satisfying the fusion rules. Hence, the topological symmetry operator provides a natural basis to relate defects described by a fusion category and conserved charges described by a ring over $\mathbb{C}$. However, this does not implement their definitions are compatible with RG flows straightforwardly, as we have demonstrated in the main text. Because RG flow is nothing but the addition of the operator to the quantum Hamiltonian, and this addition changes the role of space and time. Hence, the roles of the defects and conserved charges are deformed nontrivially through the process in general.

### v4

@@ Appendix D Verlinde operator and topological defects: $\mathbb{C}$-linear category and fusion category
where $M$ and $\overline{M}$ are the labels of the chiral or antichiral descendant fields, corresponding to the Verma module. The most fundamental point in this representation is, by definition, $P_{a}$ is commutative with the CFT Hamiltonian $H_{CFT}$, i.e. $H_{CFT},P_{a}]=0$. In other words, the above form provides representations of conserved charges. Hence, the operators form a ring over $\mathbb{C}$. Moreover, by changing the basis by a linear transformation, one can introduce the following topological symmetry operators,
@@
by identifying $\times$ as the multiplication of the linear operators. By applying the modular $S$ transformation, the operators produce defects satisfying the fusion rules. Hence, the topological symmetry operator provides a natural basis to relate defects described by a fusion category and conserved charges described by a ring over $\mathbb{C}$. However, this does not imply that their definitions are compatible with RG flows straightforwardly, as we have demonstrated in the main text. Because RG flow is nothing but the addition of the operator to the quantum Hamiltonian, and this addition changes the role of space and time. Hence, the roles of the defects and conserved charges are deformed nontrivially through the process in general.